### v2

# Revealing Human Internal Attention Patterns from Gameplay Analysis for Reinforcement Learning

Abstract
This study introduces a novel method for revealing human internal attention patterns from gameplay data alone, leveraging offline attention techniques from reinforcement learning (RL). We propose contextualized, task-relevant (CTR) attention networks, which generate attention maps from both human and RL agent gameplay in Atari environments. To evaluate whether the human CTR maps reveal internal attention, we validate our model by quantitative and qualitative comparison to the agent maps as well as to a temporally integrated overt attention (TIOA) model based on human eye-tracking data. Our results show that human CTR maps are more sparse than the agent ones and align better with the TIOA maps. Following a qualitative visual comparison we conclude that they likely capture patterns of internal attention. As a further application, we use these maps to guide RL agents, finding that human internal attention-guided agents achieve slightly improved and more stable learning compared to baselines. This work advances the understanding of human-agent attention differences and provides a new approach for extracting and validating internal attention from behavioral data.
keywords:
Visual Attention, Explainable Artificial Intelligence, Cognitive Modeling, Reinforcement Learning
1 Introduction
Figure 1: The goal of this study is to reveal human internal attention patterns solely from gameplay analysis: An attention model is proposed that extracts contextually task-relevant features from human and reinforcement learning (RL) agent gameplay. A temporally integrated gaze prediction model is used as a point of comparison to confirm that the attention model successfully reveals human internal attention.
The frame problem [1] represents an ongoing challenge in artificial intelligence (AI) research: How does an agent determine what information is relevant in complex, ever-changing environments? Understanding how humans allocate attention in task-based scenarios can provide valuable insights into this problem to eventually design more intelligent, autonomous agents. By integrating research on human attention with visual attention-based reinforcement learning (RL), we can gain a deeper understanding of attention’s role in human task learning while also improving the design of intelligent, autonomous agents.
As visual attention methods become more common in deep learning [2], several recent studies have explored incorporating attention mechanisms including human attention information into RL. These methods aim to either make an agent’s actions more interpretable, contributing to explainable AI, or improve learning efficiency and final performance. The attention approaches can be classified as offline-trainable and online-trainable. In offline-trainable attention methods such as [3, 4, 5, 6, 7, 8, 9, 10], an attention or saliency map is derived in a self-supervised way—either from the state (i.e., visual input) or from state-action-reward transitions (i.e., experience replay). This map is then used for pre-processing the input image or features by weighting or selection before processing in the policy network, or can even be applied after training for interpretation purposes too. In online methods such as [11, 12, 13, 14], attention is integrated directly into the learning process, for example, by being determined bymadditional attentive actions assigned to the agent, as in active vision reinforcement learning [11]. These attention mechanisms have been widely used in visual deep RL for Atari games but have also been applied to robotic navigation [15] and manipulation tasks [16].
Apart from applying an agent’s own attention map, agents can also be augmented with human attention information. This has been shown to improve learning performance in imitation learning (IL), including behavioral cloning [17] and inverse RL [18, 19]. Human attention information can be obtained through recording eye-tracking data during gameplay [17], and predicted using gaze prediction models [20, 21]. There is also research on task-oriented human gaze and attention prediction. In [20], a top-down gaze prediction model is built using a dynamic Bayesian network to predict human gaze positions during gameplay. Yang et al. [21] used inverse reinforcement learning to model task-oriented attention in visual search paths on images. However, to fully understand human decision making in task-based scenarios, actual or predicted gaze data alone is not sufficient. This is because gaze data only reveals overt attention [22], i.e., the attention visible to an observer. Covert attention is the mental focus on something without necessarily looking at it—such as an object in the visual periphery or in memory. Because it involves no observable movements, it can not be measured by eye tracking and is challenging to model [20, 23, 24].
Guo et al. [23] have done an in-depth comparison of human gaze prediction with RL agents’ attention (saliency map), leading to the insight that humans attend to fewer, less long-term objects. Also, agents tended to fail in situations where their attention strongly differed from the human attention. This study points out the limitations of using gaze data (overt attention) only to study human decision making and to build better human attention-augmented agents. If we had access to human internal attention maps—including both task-relevant overt and covert attention-they could be used to further improve intelligent, autonomous agent design. Without the need for additional brain activity data (such as in [24]), it remains to be explored whether task-relevant internal attention can be inferred from gameplay.
In this paper, we define human internal attention as the combination of task-relevant overt attention (observable through gaze), covert attention (mental focus not accompanied by eye movements), and short-term visual memory, all corresponding to a given state. Together, these represent the full set of task-relevant objects that guide decision making. We argue that this internal attention is of particular importance for studying human decision making and for developing better human attention-augmented agents.
The two main contributions of this study are the following: (1) We derive internal attention patterns from gameplay. The approach of this study is visualized in Fig. 1. (1a) We propose a contextualized, task-relevant (CTR) attention network trained hand-in-hand with a human action prediction network, inspired by offline attention methods in reinforcement learning (similar to visual attention in behavioral cloning [25]). Applied to the gameplay (experience replay) of humans and RL agents playing Atari games, this network helps to select features relevant to the current task and context at various sparsity levels. We first validate the CTR network using generated human and agent attention maps. (1b) We compare these attention maps against a custom temporally-integrated overt attention (TIOA) network, trained on eye-tracking data from the same human Atari gameplay. The TIOA serves as a point of comparison for covert attention, accounting for visual integration but not peripheral attention. Through this comparative analysis, we assess the CTR network’s ability to reveal internal attention patterns and provide insights that go beyond previous gaze-based studies. (2) We apply the derived internal attention maps for learning, guiding RL agents and evaluating their impact on learning performance. By leveraging these maps, we aim to transfer not only actions but also underlying attentional strategies, offering a complementary approach to imitation learning that may improve efficiency, robustness, and interpretability.
In Sec. 2, firstly the experimental setup and used datasets are explained and the architecture of the CTR attention and TIOA network as well as the temporal integration of eye-tracking data explained. The evaluation is split in three sections, where in Sec. 3 the action predictor is evaluated, in Sec. 4 the revealed attention maps are analyzed quantitatively and qualitatively, and in Sec. 5 the application of attention to reinforcement learning is presented. The results are discussed and concluded in section 6.
2 Methods
Figure 2: a) Contextualized task-relevant (CTR) attention network as well as temporally-integrated overt attention (TIOA) network architecture; b) Action predictor architecture; c) Autoencoder (AE) architecture; d) Method of creating a blended feature space $\mathcal{F}_{\mathrm{B}}$ from a target and source for the action predictor.
Figure 3: Step-by-step construction of a temporally-integrated overt attention target map from gaze positions for training of the TIOA network.
2.1 Study Overview and Datasets
The study plan illustrated in Fig. 1 will be implemented using Atari gameplay data from both humans and reinforcement learning agents, including human eye-tracking information. Atari games are chosen as a proven research platform for visual attention-based reinforcement learning. Their limited exploratory freedom further ensures that humans and agents encounter similar states despite differences in performance, making them well-suited for this comparison. Six Atari games (Freeway, MsPacman, Enduro, Seaquest, SpaceInvaders, Riverraid) are chosen to cover a variety of game types, visuals, and action counts.
[…]
For each game, and for both agent and human replay memory, we train one attention model (CTR attention) and one human gaze predictor (TIOA model). Training the attention network also involves an autoencoder and an action predictor. To prevent overreliance on elements apart from the main scene, interface elements and score indicators are deactivated. The neural network models are implemented and trained in PyTorch, with replay memory randomly split into 80% training and 20% validation transitions. All models are trained for at least 300 epochs until convergence. [^2] The code is made publicly available at: https://github.com/UThenrik/revealing-attention
2.2 Contextualized, Task-Relevant (CTR) Attention Network
We propose a contextualized, task-relevant (CTR) attention network, visualized in Fig. 2(a), which predicts an attention map
$\bm{\Psi}=f_{\mathrm{CTR}}(\bm{s},\lambda,\theta)\in\mathbb{R}^{1\times 21\times 21}$ (1)
based on the current state $\bm{s}\in\mathbb{R}^{4\times 84\times 84}$ and a sparsity controlling value $\lambda\in[0,1]$, parameterized by $\theta$. The state is represented as a stack of four game frames, each being a grayscale image of size $3\times 84\times 84$, downscaled from an RGB image of size $3\times 210\times 160$ as commonly done in Atari RL research. The state is first processed through a three-stage convolutional encoder, similar to the architecture in [6], reducing spatial resolution fourfold. The encoder weights are initialized from a pre-trained autoencoder, as illustrated in Fig. 2(b) and described in the next paragraph. After encoding, the features are contextualized using four serial dilated convolution layers with increasing dilation rates, concatenated with global maximum and average pooling channels, total spatial coordinates (similar to CoordConv [28]), and the expanded attention sparsity value $\lambda$. The final attention map is generated through three serial pointwise convolution layers, outputting a single-channel map of size $1\times 21\times 21$. Leaky ReLU activations are applied between layers, with a sigmoid activation at the output to ensure values remain in the $[0,1]$ range.
For training the CTR attention network, we use a pre-trained autoencoder (AE), as illustrated in Fig. 2(b). This AE is trained for each game separately on states from both agent and human replay memory and encodes the state into a feature tensor $\bm{\mathrm{F}}\in\mathbb{R}^{32\times 21\times 21}$ with 32 feature channels and spatial dimensions matching $\bm{\Psi}$. After the first two encoder layers we apply BlurPooling [29] with stride of 2 to ensure the feature space is shift-invariant while allowing for fourfold reduction of spatial resolution.
By applying the attention map at this encoded level, we achieve a balanced mix of spatial, feature-based, and object-based attention. For example, a single feature can preserve spatial location while also encoding an object (e.g., a car’s position in Freeway) and/or a higher-level representation (e.g., the car’s driving direction). The CTR attention network is trained alongside an action predictor, whose architecture is shown in Fig. 2(c). The action predictor estimates the action for the current target state based on a feature space:
$\hat{\bm{a}}=f_{\mathrm{AP}}(\bm{\mathrm{F}},\theta)$ (2)
where $f_{\mathrm{AP}}$ denotes the action predictor network and $\bm{\mathrm{F}}$ is a feature tensor, which can be the blended feature space, an attention-masked feature space, or the raw feature space.
$\bm{\mathrm{F}}_{\mathrm{B}}=\bm{\mathrm{F}}_{\mathrm{T}}\odot\bm{\Psi}_{\mathrm{T}}+\bm{\mathrm{F}}_{\mathrm{S}}\odot\big((1-\bm{\Psi}_{\mathrm{T}})\odot(1-\bm{\Psi}_{\mathrm{S}})\big)^{\circ\frac{1}{2}}$ (3)
This blended space overlays the target feature tensor $\bm{\mathrm{F}}_{\mathrm{T}}$ and the source feature tensor $\bm{\mathrm{F}}_{\mathrm{S}}$, which is derived from a randomly sampled state. Here, $\odot$ and $\circ$ denote element-wise multiplication and exponentiation, respectively. The blending process for the feature spaces is visualized in Fig. 2(d).
The total loss for joint training of the CTR attention network and the action predictor is
$\mathcal{L}_{\mathrm{total}}=\;$ $\mathcal{L}_{\mathrm{AP,B}}+\mathcal{L}_{\mathrm{AP}}$ (4)
$+\beta\,\mathcal{L}_{\lambda,1}+(1-\beta)\,\mathcal{L}_{\lambda,2}$ (5)
where
$\mathcal{L}_{\lambda,1}$ $=\frac{1}{n_{\lambda}}\sum_{i=1}^{n_{\lambda}}\frac{|\overline{\bm{\Psi}}_{i}-\lambda_{i}|}{|\lambda_{i}|+\epsilon}$ (6)
$\mathcal{L}_{\lambda,2}$ $=\mathcal{L}_{\lambda,1}\cdot\mathbb{I}(\overline{\bm{\Psi}}_{i}>\lambda_{i})$ (7)
and $\beta=\max(0,1-\frac{j_{\mathrm{epoch}}}{100})$ controls loss weighting over training epochs. Here, $n_{\lambda}$ is the number of sampled $\lambda$ values per batch (we use $n_{\lambda}=6$), $\overline{\bm{\Psi}}_{i}$ is the mean of the attention map for sample $i$, $\lambda_{i}$ is the corresponding sparsity target, $\epsilon$ is a small constant for numerical stability, and $\mathbb{I}(\cdot)$ is the indicator function. We use the cross-entropy
$\mathrm{CE}(\bm{a},\hat{\bm{a}})=-\frac{1}{n_{\mathrm{b}}}\sum_{j=1}^{n_{\mathrm{b}}}\sum_{i=1}^{n_{\mathrm{a}}}a_{i,j}\log(\hat{a}_{i,j})$ (8)
where $\bm{a}$ is the true action (one-hot encoded), $\hat{\bm{a}}$ is the predicted action probability, $n_{\mathrm{b}}$ is the batch size, and $n_{\mathrm{a}}$ is the number of actions. The action prediction losses are then given by
$\mathcal{L}_{\mathrm{AP,B}}$ $=\mathrm{CE}(\bm{a},\hat{\bm{a}}^{\mathrm{B}})$ (9)
[…]
where $\hat{\bm{a}}^{\mathrm{B}}$ and $\hat{\bm{a}}$ are the predicted action probabilities from the blended and full feature spaces, respectively. The attention maps used for the blended feature space action prediction loss $\mathcal{L}_{\mathrm{AP,B}}$ are the same as those used for the sparsity losses $\mathcal{L}_{\lambda,1}$ and $\mathcal{L}_{\lambda,2}$, which is why concurrent training of the CTR attention network and action predictor is required.
The sparsity-controlling factor $\lambda$ restricts the average activation of the attention map, specifying the maximum ratio of features that can be active. The loss term $\mathcal{L}_{\lambda,2}$ enforces this. While we assume that more features may improve action prediction by providing more information, we want to avoid unnecessary activations. The auxiliary loss $\mathcal{L}_{\lambda,1}$ also controlling minimum sparsity in early training is introducted for training stabilization.
The total loss (4) combines cross-entropy terms, ensuring the action predictor selects the correct action, with a sparsity regularization term that keeps the attention map controllable. The blended feature space encourages the CTR attention network to focus on contextually task-relevant features from the target and suppress irrelevant features from the source, driving $\bm{\Psi}$ values toward 0 or $1$. As a result, the CTR attention network learns to spatially extract features relevant to the player’s decision in each gameplay context, adapting flexibly to different sparsity levels specified by $\lambda$. For example, the network is trained such that, if we input $\lambda=0.05$, it activates on average at most 5% of the features.
The action predictor is trained using a learning rate of $1\times 10^{-3}$, while the CTR attention network is trained with half that rate, i.e., $5\times 10^{-4}$, which encourages stable joint optimization. After training, the CTR attention models can be used independently of the autoencoder and action predictor.
It is noted that the CTR attention network can be used plug-and-play with the self-supervised attention network from [6], which has been shown to improve the rate of convergence and performance of deep RL methods in Atari games, while additionally offering task-relevant and contextualized feature selection. This will therefore be investigated in Sec. 5. It could also be trained using inverse dynamics prediction, i.e. predicting an action $a_{k}$ from the current state $\bm{s}_{k}$ and the next $\bm{s}_{k+1}$. This approach would extract all features that are evidence of an action and could be more robust for applications in RL. Forward action prediction, equivalent to behavioral cloning, is used in this study as it leads the attention map to extract features that likely caused the player’s decision.
2.3 Temporally-Integrated Overt Attention (TIOA) Network
In order to analyze whether the CTR network can reveal human internal attention, a point of comparison is needed. Rather than using only the instantaneous gaze positions (eye-tracking data) from the Atari-HEAD dataset, an integration of recent gaze positions is preferable, as it more closely mimics the temporal integration of visual sensory data in humans [30]. This is relevant for comparison in accordance to our definition in Sec. 1 where human internal attention encompasses short-term visual memory as well. To further enhance generalization to unseen states (and those of the agent), we propose a temporally-integrated overt attention (TIOA) network. It features the same base architecture as the CTR attention network in Fig. 2(a) and therefore predicts temporally-integrated overt attention on the fourfold resolution-reduced feature space. Its output shall denote a probability distribution and therefore a softmax function is used after the final layer.
[…]
$w(\tau)=1.1-10^{\frac{\tau}{c}-1}$ (11)
[…]
$\mathcal{L}_{\mathrm{TIOA}}=\frac{1}{n_{\mathrm{b}}}\sum_{j=1}^{n_{\mathrm{b}}}\sum_{i}^{n_{\mathrm{f}}}\hat{\bm{\Gamma}}_{i,j}\log\frac{\hat{\bm{\Gamma}}_{i,j}}{\bm{\Gamma}_{i,j}^{*}}$ (12)
between the temporally integrated gaze target map $\bm{\Gamma}^{*}$ and its prediction $\hat{\bm{\Gamma}}$ with the number of spatial features ${n_{\mathrm{f}}}=441$.
3 Evaluation I: Action Prediction under Attention Masking
Figure 4: Action predictor accuracies over unseen attention-masked states for human (CTR and TIOA) and agent (only CTR) gameplay at different attention activation rates/ sparsities.
In this section, we evaluate the action prediction accuracy for both human and agent gameplay under full observation and under attention masking. For the human action prediction, we assess attention masking using both CTR and TIOA attention maps of varying sparsity that is controlled by the sparsity value $\lambda$. For the agent, we evaluate attention masking using only the CTR attention maps. The action predictors used here are trained directly on the attention-masked states produced by the trained CTR and TIOA networks. Fig. 4 shows the resulting action predictor accuracies
$p_{\mathrm{A}}=\frac{1}{n_{\mathrm{s}}}\sum_{j=1}^{n_{\mathrm{s}}}\mathbb{I}\Big(\arg\max_{i}\hat{a}_{i,j}=\arg\max_{i}a_{i,j}\Big)$ (13)
over all unseen states $n_{\mathrm{s}}$ for human and agent gameplay, with activation rates $r_{\Psi}$ and $r_{\Gamma}$ evaluated at $0.02$, $0.04$, $0.08$, $0.16$, and $1.0$. For the CTR attention maps, the activation rate $r_{\Psi}$ is achieved by controlling $\lambda$ so that as the average activation in the unseen states results to:
$r_{\Psi}\approx\frac{1}{n_{\mathrm{s}}}\sum_{i=1}^{n_{\mathrm{s}}}\overline{\Psi}_{i}$ (14)
For TIOA, which produces a probability map, we binarize the attention map by selecting the top $r_{\Gamma}$ fraction of features (i.e., those with the highest logits), setting them to $1$ and the rest to 0:
$\hat{\bm{\Gamma}}=\mathbb{I}\left(\hat{\Gamma}\geq\tau_{k}\right)$ (15)
where $\tau_{k}$ is the threshold such that $r_{\Gamma}$ fraction of features are activated.
We observe that accuracies for the human actions are higher than in the original Atari-HEAD study [17], likely as we are predicting based on frame-stacked states.
In total, four main findings can be observed. Firstly, higher activation rates, i.e., more information leads to more accurate action prediction. This is the case for all games except Freeway. Secondly, the CTR attention maps are more efficient than the gaze data-based TIOA maps, achieving significantly higher accuracies with fewer activated features. This may be expected, as the CTR networks are optimized for this task. However it underlines the core motivation for this study as posed in Sec. 1, that the full internal attention including covert attention is required to better describe human decision making. The dashed lines in Fig. 4 indicate the minimum activation rate where 95% of the maximum action prediction accuracy is reached. Thirdly, human actions can be accurately predicted with less features than agent actions. This agrees with the finding of Guo et al. [23] which was based only on gaze data. Humans focus on less objects than agents trained for long-term future reward, even if we include potential objects of covert attention in the periphery beyond the gaze. Lastly, despite the expected greater gameplay variety in the human replay memory (due to the involvement of four different human subjects), the prediction accuracies are similar to those of a replay memory of a single trained agent.
The findings suggest that the CTR network correctly identifies contextually task-relevant features. Whether these maps plausibly resemble human internal attention patterns will be further analyzed quantitatively and qualitatively in Sec. 4.
4 Evaluation II: Analysis Revealed Attention Maps
4.1 Visual Analysis of Attention Maps
Figure 5: Binarized attention map comparison: Example scenes (unseen states) with binarized TIOA (upper) and CTR attention map (lower) overlay. CTR human attention maps more closely resemble the TIOA than the agent attention. They likely represent human interal attention patterns as they include parts of the TIOA map, objects nearby, and highly task-relevant objects (see annotations).
Figure 6: Smooth attention map comparison: Example scenes (unseen states) with TIOA and Human CTR attention map that have been evaluated at multiple sparsity levels, smoothed and averaged. The MsPacman frame has been converted to grayscale for better visualization.
In Fig. 5, a scene is presented for each game with overlays of the TIOA and CTR maps. The selected scenes are intended to represent typical characteristics of the TIOA and CTR maps. Animations are provided in the supplemental materials. To adjust for differences between human and agent attention differences identified in Sec. 3 we base the chosen activation rates $r_{\Psi}$ off the 95% of action prediction accuracy threshold from 4, with a lower limit of $0.04$ and an upper limit of $0.16$ resulting in the values shown in Tab. 1. In Fig. 5 for better visualization, we use a stricter upper limit of $0.08$.
   Game    Agent    Human
   Enduro    0.16    0.097
   Freeway    0.096    0.04
   MsPacman    0.16    0.077
   Seaquest    0.143    0.04
   SpaceInvaders    0.16    0.04
   Riverraid    0.16    0.041
Table 1: CTR attention map activation rates $r_{\psi}$ for each game used in the analysis in Sec. 4.
For the visualization of the CTR attention maps in Fig. 5, the maps are binarized at a threshold of $0.5$, and then upscaled. The TIOA map probabilities are binarized at a threshold in the range $[0.005,0.01]$ depending on the game [^3] TIOA binarization thresholds are: $0.005$ (Freeway), $0.0075$ (Seaquest), $0.01$ (Enduro, MsPacman, SpaceInvaders, Riverraid).. Providing additional scenes, Fig. 6 shows the TIOA and Human CTR attention maps. The latter have been evaluated at multiple sparsity levels [^4] The target sparsity levels used for the smoothed Human CTR maps are: $0.01$, $0.02$, $0.04$, and $0.08$., smoothed with a Gaussian kernel of standard deviation $1$ and averaged. This is to make use of the controllable sparsity to create a continuous attention map with similar distribution and appearance as the (un-binarized) TIOA maps. Lastly, Fig. 7 presents overall heatmaps at the sparsities from Tab. 1 over all unseen states (validation dataset).
It will be qualitatively analyzed for each game whether the TIOA and attention maps differ, and if the CTR attention maps plausibly resemble internal attention. Again, the TIOA maps serve as a point of comparison as no true label of the internal attention is available. Generally, the TIOA maps are expected to overlap better with the human CTR attention maps but not fully.
In Enduro, the TIOA predicts gaze toward the distant road and sometimes incoming cars, while the CTR similarly focuses on the cars and road boundaries, also incorporating the avatar’s car. The human seems to attend further ahead, potentially due to slower reaction time than the agent. Given its significance in the game, the vatar’s car is likely an object of covert attention or memory, as it can be perceived both peripherally and its position can be memorized. The heatmaps all show a focus towards the center.
In Freeway, the agent consistently attends to most cars and the player’s avatar (a chicken crossing the road). In contrast, human CTR attention alternates between two distinct modes: when walking, the focus is mostly on the avatar (seen in 6), and when walking, attention shifts toward the road ahead and cars (seen in 5). The 2 modes can also be seen in quantiative analysis in the following Sec. 4.2. The pattern aligns with human gaze behavior, where attention is primarily directed at the path ahead and objects near it. The human generally focuses stronger on the line the avatar is walking than the agent. This is predicted by the CTR network and can be seen clearly in the heatmap as well.
For MsPacman and Seaquest, the TIOA predicts human gaze primarily around the avatar character and a few opponents or collectibles. The human CTR attention map likely reflects a pattern of interal attention, especially in MsPacman, where the similarities are strongest. In the shown scene, the CTR attention map focuses on a relevant nearby ghost which is also looked at by the human avatar, but rather at the edge of focus area. The agent’s CTR map, on the other hand, extends attention to opponents and smaller items farther away. The agent’s attention is stronger along the horizontal line along the avatar’s vertical position. Notably, heatmaps show that both the human CTR and TIOA maps attend to various positions in the game environment.
In Space Invaders, the TIOA predicts gaze primarily on the attacking aliens rather than the avatar cannon. The agent CTR map, in contrast, distributes attention across both the avatar and most aliens. The human CTR map also focuses on the shields and slightly prioritizes the aliens immediately above the avatar while ignoring others, reflecting a more selective attention strategy.
In Riverraid, the TIOA predicts attention toward incoming targets and fuel fields ahead of the avatar’s plane, primarily in the middle and upper third of the screen, as well as occasionally on the plane itself. The human CTR map focuses on most dynamic objects, without clearly distinguishing between incoming opponents and purely aesthetic elements. However, in Fig. 6 the attention on the aesthetic elements is much lower. The agent’s attention map also includes the river’s edge, which marks an obstacle. Riverraid is perhaps the most complex of the six games analyzed, as it features a constantly changing background and scenarios. The limited dataset size likely prevents the CTR network from extracting meaningful patterns from the human gameplay for this game and the resulting attention maps may not fully reflect internal attention patterns.
Figure 7: Heatmaps of TIOA and CTR attention maps over unseen states.
4.2 Similarity Analysis and Histograms of Attention Maps
Figure 8: Comparison for CTR agent and human attention map activation rates and alignment score toward the TIOA maps. CTR human attention maps have lower activation rates (higher sparsity) than the CTR agent maps and in average higher alignment scores (higher similarity) to the TIOA.
An analysis of the characteristics of the CTR attention maps is performed. The CTR attention maps are evaluated for a specific attention activation rate as listed in Tab. 1. $1024$ states are randomly sampled half-half from the human and agent validation datasets. Fig. 8 shows histograms for activation rate $r_{\Psi}=\bm{\bar{\Psi}}$ of the CTR attention maps. This corresponds to the ratio of features attended to among all features. We observe that the CTR activation rates are generally spread evenly around the targeted average one, where the agent ones feature a higher tail towards the higher values of $r_{\Psi}$.
Further, Fig. 8 shows alignment scores
$a_{\Psi}=\frac{1}{\overline{\bm{\Psi}}}\sum_{i}^{n_{\mathrm{f}}}\bm{\Psi}_{i}\odot\hat{\bm{\Gamma}}_{i}$ (16)
computed between CTR attention maps $\bm{\Psi}$ and the continuous TIOA gaze prediction map $\hat{\bm{\Gamma}}$. Here, $\overline{\bm{\Psi}}$ denotes the mean value of $\bm{\Psi}$, normalizing the score such that $a_{\Psi}=1$ indicates a coverage of probability mass equivalent to random attention. Values $a_{\Psi}>1$ indicate better-than-random alignment, with higher values therefore reflecting stronger similarity with TIOA predictions. We chose $a_{\Psi}$ over binary cross-entropy (BCE) as a tool for comparison because it provides an interpretable, activation-rate-invariant measure of relative concentration. In contrast, the expected BCE for a fixed predictor depends on the overall activation rate of $\Psi$, with higher activation generally leading to higher BCE (thus unfairly favouring the more sparse human CTR maps).
We observe that both human and agent CTR attention maps outperform the baseline of $1.0$. For all games, the human CTR maps reach a higher alignment score averaging at $3.26$ compared to the agent CTR maps, which reach $2.49$. The alignment scores vary across games, with the highest human-agent gap in Freeway and MsPacman, indicating that human attention patterns in these games are more closely aligned with TIOA predictions. Space Invaders and Riverraid show the smallest differences. For Freeway, we observe a long tail end of the histogram, this reflects the mode where the human is predicted to focus on the road ahead as described in Sec. 3. Interestingly, the alignment score per bin also spikes here as well, underlying that this pattern exhibits consistency with the TIOA predictions. Overall, these results demonstrate that human CTR attention maps consistently exhibit a strong similarity to the TIOA maps across all games, indicating that the CTR network can capture human internal attention patterns.
5 Evaluation III: Human Internal Attention-Guided Reinforcement Learning
Figure 9: Comparison of learning curves of the Dueling DDQN agent with PER for the plain agent, the human internal attention-guided agent (CTR), and the inverted attention agent. Three runs are averaged per agent type, after performing a rolling average of 200 steps per run. Shaded area indicates standard deviation. Overall, CTR-masked agents slightly outperform plain ones, showing stable learning.
Achieving human scores in Atari games is a challenging benchmark for RL agents, especially under limited training steps. Even algorithms such as Rainbow Deep Q-Learning (DQN) [32] that combine several improvements over normal DQN fail to reach mean human-normalized performance within 10 million environment steps, while human-outperforming methods like Agent57 [33] are highly tailored for Atari. Guo et al. [23] found that RL agents (here, using Proximal Policy Optimization, PPO), when they fail, are more likely to neglect regions attended to by humans. We investigate whether the human internal attention maps can improve learning performance, by making RL agents focus on objects that are task-relevant for humans. This represents an alternative approach to imitation learning, as it aims to transfer not just actions but the underlying attentional strategies derived from those actions.
As a strong baseline, we use DQN with the modifications of double DQN (DDQN) [34], dueling DQN [35], and proportional prioritized experience replay (PER) [36]. Dueling DQN improves learning stability by separating value and advantage streams, double DQN mitigates overestimation bias by using a separate target network, and PER allows the agent to learn more effectively from rare but significant events. These enhancements are implemented using a custom-modified version of the stable baselines 3 (SB3) [37] DQN implementation. The policy network receives the feature representation $\mathcal{F}_{\mathrm{p}}$ as input. We evaluate three agent variants: (1) a plain agent using features from a pre-trained, frozen autoencoder ($\mathcal{F}_{\mathrm{p}}=\mathcal{F}$), (2) a human internal attention-guided agent, where features are masked by the attention map $\bm{\Psi}$ to focus on task-relevant objects ($\mathcal{F}_{\mathrm{p}}=\bm{\Psi}\odot\mathcal{F}$), and (3) an inverted attention agent, where features are masked by the inverted CTR map ($\mathcal{F}_{\mathrm{p}}=(1-\bm{\Psi})\odot\mathcal{F}$). The comparison to the inverted map is done similarly to [38], providing a control baseline to evaluate the usefulness of the attention maps, especially since selective attention does not always guarantee performance gains [39]. For each agent (1-3), we perform three runs of different seeds. For all experiments, we use $16\%$ average activation rate in the CTR attention maps (meaning $84\%$ average activation for the inverted maps).
We set the Dueling DDQN and PER hyperparameters to standard values commonly used in the literature. Training is performed for 2.5 million steps (with frameskip 4, corresponding to 10 million frames), using a buffer size of 500,000, batch size of 32, and a learning rate of $2.5\times 10^{-4}$. The agent starts learning after 25,000 steps, updates every 4 steps, and uses a target network update interval of 10,000. The discount factor is set to $0.99$, and exploration follows a linear schedule from $1.0$ to $0.1$ over the first 10% of training.
Fig. 9 shows the learning curves for the three agent variants, denoting the reward obtained during the actual training episodes. The human internal attention-guided agent outperforms the plain agent for Enduro, Freeway and MsPacman, achieves similar performance for SpaceInvaders and slightly worse for Seaquest and Riverraid. The inverted attention agent performs worse in most games, especially in Seaquest where almost no learning progress is made. Unintuitively, its performance in Freeway is similar to the plain agent. The CTR-masked agents also appear to learn quite stably, especially in Enduro and MsPacman.
Figure 10: Comparison of episodic scores of the Dueling DDQN agent with PER for the plain agent, the human internal attention-guided agent (CTR), and the inverted attention agent, including the random agent score. Three runs are averaged per agent type. Shaded area indicates standard deviation. Overall, CTR-masked agents consistently achieve higher episodic scores compared to plain agents.
Fig. 10 shows the average episodic scores for the three agent variants after training. Here, the agents are evaluated under an exploration rate of $\epsilon=0.01$ for 150 episodes (50 per agent of different seeds), meaning they exploit their learned policies while still choosing a random action in $1\%$ of the cases. An overall average relative score is determined by
$\bar{R}^{(m)}=\frac{1}{n_{\mathrm{g}}}\sum_{g=1}^{n_{\mathrm{g}}}\frac{S_{g}^{(m)}-S_{g}^{(\mathrm{rand})}}{S_{g}^{(\mathrm{plain})}-S_{g}^{(\mathrm{rand})}}$ (17)
is determined using random scores [^5] Random agent scores: Enduro: 0.0, Freeway: 0.0, MsPacman: 307.3, Seaquest: 68.4, SpaceInvaders: 148.0, Riverraid: 1338.5, taken from [33] and plain scores as the baselines. The CTR-masked agent achieves highest episodic scores for MsPacman, Riverraid, and SpaceInvaders. It clearly outperforms the inverted attention agent in all games except Freeway, where all three agent types perform similarly. For the average relative score, the CTR-masked agents achieve $1.10$ whereas the inverted attention agents underperform with a score of $0.68$. The human internal attention-guided agents therefore reach $10\%$ higher scores than the baseline Dueling DDQN agent with PER. This shows that the derived attention maps can slightly improve RL agents performance, limiting the policy network’s available feature space to the $16\%$ predicted to be most relevant for humans.
6 Conclusions
Our study presents a novel method to derive human internal attention patterns from gameplay data alone, leveraging offline attention methods from reinforcement learning (RL). By introducing the contextualized, task-relevant (CTR) attention network, we generate attention maps from human and agent gameplay, compared against a temporally-integrated overt attention (TIOA) model based on eye-tracking data. After qualitative and quantitative evaluation, these maps are applied for attention-guided reinforcement learning.
Evaluation across six Atari games demonstrates that human CTR attention maps exhibit higher alignment towrad the TIOA maps compared to those generated for agents. Overall, the human CTR maps capture key regions of the TIOA maps, as well as adjacent objects and highly task-relevant features, providing evidence that the CTR network reveals internal human attention patterns. Applied to RL, CTR-masked agents achieved 10% higher scores than baseline using only 16% of the feature space, demonstrating improved learning efficiency and performance, offering an alternative to imitation learning approaches.
Limitations include the need for larger human gameplay datasets with eye-tracking to improve accuracy and reduce noise in the CTR maps. Larger datasets could also allow studying differences across subjects. Methods could further be extended to datasets from autonomous driving or robotics.
Our results represent a first step toward predicting human internal attention from interactions alone, without relying on additional data such as brain activity. This approach can deepen understanding of human-agent attention differences and contribute to human-inspired autonomous agents. Future work could explore more advanced strategies for human internal attention-guided learning, aiding the development of explainable, interpretable, and robust autonomous systems.
[…]
Supplementary information
We provide supplementary animations of the CTR and TIOA attention maps, namely the binarized maps shown in Fig. 5, the smoothed maps shown in Fig. 6, as well as the maps binarized at average activation of 16% as used in the RL experiments in Sec. 5.
Declarations
The authors did not receive support from any organization for the submitted work.

### v3

@@
-Revealing Human Internal Attention Patterns from Gameplay Analysis for Reinforcement Learning
+Revealing Human Attention Patterns from Gameplay Analysis for Reinforcement Learning
 Abstract
-This study introduces a novel method for revealing human internal attention patterns from gameplay data alone, leveraging offline attention techniques from reinforcement learning (RL).
+This study introduces a novel method for revealing human internal attention patterns (decision-relevant attention) from gameplay data alone, leveraging offline attention techniques from reinforcement learning (RL).
 We propose contextualized, task-relevant (CTR) attention networks, which generate attention maps from both human and RL agent gameplay in Atari environments.
-To evaluate whether the human CTR maps reveal internal attention, we validate our model by quantitative and qualitative comparison to the agent maps as well as to a temporally integrated overt attention (TIOA) model based on human eye-tracking data.
+To evaluate whether the human CTR maps reveal internal attention patterns, we validate our model by quantitative and qualitative comparison to the agent maps as well as to a temporally integrated overt attention (TIOA) model based on human eye-tracking data.
 Our results show that human CTR maps are more sparse than the agent ones and align better with the TIOA maps.
 Following a qualitative visual comparison we conclude that they likely capture patterns of internal attention.
-As a further application, we use these maps to guide RL agents, finding that human internal attention-guided agents achieve slightly improved and more stable learning compared to baselines.
-This work advances the understanding of human-agent attention differences and provides a new approach for extracting and validating internal attention from behavioral data.
+As a further application, we use these maps to guide RL agents, finding that human attention-guided agents achieve slightly improved and more stable learning compared to baselines, and significantly outperform TIOA-based agents.
+This work advances the understanding of human-agent attention differences and provides a new approach for extracting and validating internal attention patterns from behavioral data.
 keywords:
@@
 1 Introduction
-Figure 1: The goal of this study is to reveal human internal attention patterns solely from gameplay analysis: An attention model is proposed that extracts contextually task-relevant features from human and reinforcement learning (RL) agent gameplay.
-A temporally integrated gaze prediction model is used as a point of comparison to confirm that the attention model successfully reveals human internal attention.
-The frame problem [1] represents an ongoing challenge in artificial intelligence (AI) research: How does an agent determine what information is relevant in complex, ever-changing environments?
+Figure 1: The goal of this study is to reveal human internal attention patterns (decision-relevant attention) solely from gameplay analysis: An attention model is proposed that extracts contextually task-relevant features from human and reinforcement learning (RL) agent gameplay.
+A temporally integrated gaze prediction model is used as a point of comparison to confirm that the attention model successfully reveals human internal attention patterns.
+The frame problem [mccarthy1981some] represents an ongoing challenge in artificial intelligence (AI) research: How does an agent determine what information is relevant in complex, ever-changing environments?
 Understanding how humans allocate attention in task-based scenarios can provide valuable insights into this problem to eventually design more intelligent, autonomous agents.
 By integrating research on human attention with visual attention-based reinforcement learning (RL), we can gain a deeper understanding of attention’s role in human task learning while also improving the design of intelligent, autonomous agents.
-As visual attention methods become more common in deep learning [2], several recent studies have explored incorporating attention mechanisms including human attention information into RL.
+As visual attention methods become more common in deep learning [hassanin2024visual], several recent studies have explored incorporating attention mechanisms including human attention information into RL.
 These methods aim to either make an agent’s actions more interpretable, contributing to explainable AI, or improve learning efficiency and final performance.
 The attention approaches can be classified as offline-trainable and online-trainable.
-In offline-trainable attention methods such as [3, 4, 5, 6, 7, 8, 9, 10], an attention or saliency map is derived in a self-supervised way—either from the state (i.e., visual input) or from state-action-reward transitions (i.e., experience replay).
+In offline-trainable attention methods such as [greydanus2018visualizing, manchin2019reinforcement, gopalakrishnan2020unsupervised, wu2021self, shi2020self, nikulin2019free, ma2025don, chen2024focus], an attention or saliency map is derived in a self-supervised way—either from the state (i.e., visual input) or from state-action-reward transitions (i.e., experience replay).
 This map is then used for pre-processing the input image or features by weighting or selection before processing in the policy network, or can even be applied after training for interpretation purposes too.
-In online methods such as [11, 12, 13, 14], attention is integrated directly into the learning process, for example, by being determined bymadditional attentive actions assigned to the agent, as in active vision reinforcement learning [11].
-These attention mechanisms have been widely used in visual deep RL for Atari games but have also been applied to robotic navigation [15] and manipulation tasks [16].
+In online methods such as [shang2023active, mnih2014recurrent, kulkarni2019unsupervised, mott2019towards], attention is integrated directly into the learning process, for example, by additional attentive actions assigned to the agent, as in active vision reinforcement learning [shang2023active].
+These attention mechanisms have been widely used in visual deep RL for Atari games but have also been applied to robotic navigation [salter2021attention] and manipulation tasks [james2022coarse].
 Apart from applying an agent’s own attention map, agents can also be augmented with human attention information.
-This has been shown to improve learning performance in imitation learning (IL), including behavioral cloning [17] and inverse RL [18, 19].
-Human attention information can be obtained through recording eye-tracking data during gameplay [17], and predicted using gaze prediction models [20, 21].
+Zhang et al.
+[zhang2019leveraging] provide a comprehensive survey of methods for leveraging human guidance in deep RL, categorizing attention-based learning as one approach for agents to learn from humans.
+This has been shown to improve learning performance in imitation learning (IL), including behavioral cloning [zhang2020atari] and inverse RL [zhang2018agil, saran2021efficiently].
+Human attention information can be obtained through recording eye-tracking data during gameplay [zhang2020atari], and predicted using gaze prediction models [borji2014what, yang2020predicting].
+However, as Zhang et al.
+[zhang2019leveraging] note, existing work in this area has primarily focused on using gaze (overt attention) as the sole attention signal.
 There is also research on task-oriented human gaze and attention prediction.
-In [20], a top-down gaze prediction model is built using a dynamic Bayesian network to predict human gaze positions during gameplay.
+In [borji2014what], a top-down gaze prediction model is built using a dynamic Bayesian network to predict human gaze positions during gameplay.
 Yang et al.
-[21] used inverse reinforcement learning to model task-oriented attention in visual search paths on images.
+[yang2020predicting] used inverse reinforcement learning to model task-oriented attention in visual search paths on images.
 However, to fully understand human decision making in task-based scenarios, actual or predicted gaze data alone is not sufficient.
-This is because gaze data only reveals overt attention [22], i.e., the attention visible to an observer.
+This is because gaze data only reveals overt attention [posner1980orienting], i.e., the attention visible to an observer.
 Covert attention is the mental focus on something without necessarily looking at it—such as an object in the visual periphery or in memory.
-Because it involves no observable movements, it can not be measured by eye tracking and is challenging to model [20, 23, 24].
+Because it involves no observable movements, it can not be measured by eye tracking and is challenging to model [borji2014what, guo2021machine, leong2017dynamic].
 Guo et al.
-[23] have done an in-depth comparison of human gaze prediction with RL agents’ attention (saliency map), leading to the insight that humans attend to fewer, less long-term objects.
-Also, agents tended to fail in situations where their attention strongly differed from the human attention.
+[guo2021machine] conducted an in-depth comparison of human gaze prediction with RL agents’ perturbation-based attention (saliency map), showing that agent attention becomes more similar to human gaze when the discount factor is reduced from the typical 0.99 (long-term reward focus), and that agents fail more often when their attention diverges from human gaze.
+However, their analysis relied exclusively on gaze (overt attention) and therefore did not account for covert attention or short-term memory.
 This study points out the limitations of using gaze data (overt attention) only to study human decision making and to build better human attention-augmented agents.
-If we had access to human internal attention maps—including both task-relevant overt and covert attention-they could be used to further improve intelligent, autonomous agent design.
-Without the need for additional brain activity data (such as in [24]), it remains to be explored whether task-relevant internal attention can be inferred from gameplay.
-In this paper, we define human internal attention as the combination of task-relevant overt attention (observable through gaze), covert attention (mental focus not accompanied by eye movements), and short-term visual memory, all corresponding to a given state.
-Together, these represent the full set of task-relevant objects that guide decision making.
-We argue that this internal attention is of particular importance for studying human decision making and for developing better human attention-augmented agents.
+If we had access to human attention patterns—including both decision-relevant overt and covert attention-they could be used to further improve intelligent, autonomous agent design.
+Without the need for additional brain activity data (such as in [leong2017dynamic]), it remains to be explored whether these internal attention patterns can be inferred from gameplay.
+In this paper, we define human internal attention patterns (or equivalently, decision-relevant attention patterns) as generalized patterns reflecting the combination of overt attention (observable through gaze), covert attention (mental focus not accompanied by eye movements), and short-term visual memory, all corresponding to a given state.
+Together, these represent the full set of features a human typically uses to make decisions.
+We use the term “internal” to distinguish this from agent-optimized task-relevant attention, as human internal attention is specific to human decision-making and independent of whether those decisions are optimal for task performance.
+We argue that these internal attention patterns are of particular importance for studying human decision making and for developing better human attention-augmented agents.
+To extract these patterns from gameplay, we couple action prediction with attention extraction.
+While similar methods have been used to explain agent policies (e.g., Chen et al.
+[chen2024focus], Shi et al.
+[shi2022temporal], Greydanus et al.
+[greydanus2018visualizing]), our focus is on deriving attention maps that explain human decision-making.
 The two main contributions of this study are the following: (1) We derive internal attention patterns from gameplay.
@@
 1.
-(1a) We propose a contextualized, task-relevant (CTR) attention network trained hand-in-hand with a human action prediction network, inspired by offline attention methods in reinforcement learning (similar to visual attention in behavioral cloning [25]).
-Applied to the gameplay (experience replay) of humans and RL agents playing Atari games, this network helps to select features relevant to the current task and context at various sparsity levels.
+(1a) We propose a contextualized, task-relevant (CTR) attention network trained hand-in-hand with an action prediction network, inspired by offline attention methods in reinforcement learning (similar to visual attention in behavioral cloning [mohandoss2019visual]).
+Applied to the gameplay (experience replay) of humans and RL agents playing Atari games, this network helps to select features relevant to the current state at various sparsity levels.
 We first validate the CTR network using generated human and agent attention maps.
@@
 Through this comparative analysis, we assess the CTR network’s ability to reveal internal attention patterns and provide insights that go beyond previous gaze-based studies.
-(2) We apply the derived internal attention maps for learning, guiding RL agents and evaluating their impact on learning performance.
+(2) We apply the derived internal attention patterns for learning, guiding RL agents and evaluating their impact on learning performance.
+We compare CTR-masked agents against TIOA-masked agents, demonstrating that CTR attention can offer additional value beyond overt attention alone for RL applications.
 By leveraging these maps, we aim to transfer not only actions but also underlying attentional strategies, offering a complementary approach to imitation learning that may improve efficiency, robustness, and interpretability.
@@
 2 Methods
-Figure 2: a) Contextualized task-relevant (CTR) attention network as well as temporally-integrated overt attention (TIOA) network architecture; b) Action predictor architecture; c) Autoencoder (AE) architecture; d) Method of creating a blended feature space $\mathcal{F}_{\mathrm{B}}$ from a target and source for the action predictor.
+e)Refer to caption
+Figure 2: a) Contextualized task-relevant (CTR) attention network as well as temporally-integrated overt attention (TIOA) network architecture; b) Autoencoder (AE) architecture; c) Action predictor architecture; d) Method of creating a blended feature space $\mathcal{F}_{\mathrm{B}}$ from a target and source for the action predictor; e) Exmple attention map and ablation study: Example frame and attention map (left), and log10 histogram of activations (right) over different sparsity control factors $\lambda$ for the human CTR attention network with and without feature blending.
+Feature blending drives the activation distribution toward binarization (values near 0 or 1).
 Figure 3: Step-by-step construction of a temporally-integrated overt attention target map from gaze positions for training of the TIOA network.
@@
 […]
+Although four subjects were recorded for each game, the available data per individual is deemed insufficient for robust subject-specific analysis.
+We therefore aggregate the gameplay data from all subjects to focus on more general differences between human and agent attention patterns.
+Consequently, an analysis of individual differences among subjects was not performed in this study.
 For each game, and for both agent and human replay memory, we train one attention model (CTR attention) and one human gaze predictor (TIOA model).
@@
 The state is represented as a stack of four game frames, each being a grayscale image of size $3\times 84\times 84$, downscaled from an RGB image of size $3\times 210\times 160$ as commonly done in Atari RL research.
-The state is first processed through a three-stage convolutional encoder, similar to the architecture in [6], reducing spatial resolution fourfold.
+The state is first processed through a three-stage convolutional encoder, similar to the architecture in [wu2021self], reducing spatial resolution fourfold.
 The encoder weights are initialized from a pre-trained autoencoder, as illustrated in Fig.
 2(b) and described in the next paragraph.
-After encoding, the features are contextualized using four serial dilated convolution layers with increasing dilation rates, concatenated with global maximum and average pooling channels, total spatial coordinates (similar to CoordConv [28]), and the expanded attention sparsity value $\lambda$.
+After encoding, the features are contextualized using four serial dilated convolution layers with increasing dilation rates, concatenated with global maximum and average pooling channels, total spatial coordinates (similar to CoordConv [liu2018intriguing]), and the expanded attention sparsity value $\lambda$.
 The final attention map is generated through three serial pointwise convolution layers, outputting a single-channel map of size $1\times 21\times 21$.
@@
 This AE is trained for each game separately on states from both agent and human replay memory and encodes the state into a feature tensor $\bm{\mathrm{F}}\in\mathbb{R}^{32\times 21\times 21}$ with 32 feature channels and spatial dimensions matching $\bm{\Psi}$.
-After the first two encoder layers we apply BlurPooling [29] with stride of 2 to ensure the feature space is shift-invariant while allowing for fourfold reduction of spatial resolution.
+After the first two encoder layers we apply BlurPooling [pmlr-v97-zhang19a] with stride of 2 to ensure the feature space is shift-invariant while allowing for fourfold reduction of spatial resolution.
 By applying the attention map at this encoded level, we achieve a balanced mix of spatial, feature-based, and object-based attention.
@@
 2(c).
+Crucially, we architecturally separate the attention network from the action predictor: the CTR network produces masks that are applied before the action predictor.
+This design ensures that extracted attention maps reflect genuine decision-relevant features rather than being artifacts of the action prediction optimization process itself.
 The action predictor estimates the action for the current target state based on a feature space:
@@
 2(d).
+This feature blending mechanism drives the CTR attention network toward binarized attention maps.
+By requiring the network to distinguish between target and source features while maintaining action prediction accuracy, attention values are encouraged to converge toward either 0 (irrelevant) or 1 (relevant).
+An ablation study in Fig.
+2(e) shows that without feature blending, the network produces less binarized activation distributions.
 The total loss for joint training of the CTR attention network and the action predictor is
@@
 The auxiliary loss $\mathcal{L}_{\lambda,1}$ also controlling minimum sparsity in early training is introducted for training stabilization.
+The sparsity factor $\lambda$ is implemented as a controllable network input, enabling post-training adjustment across multiple sparsity levels essential for our comparative analysis in Sec.
+3.
+Without regularization, attention maps converge to full activation to maximize information, defeating the purpose of selective attention.
 The total loss (4) combines cross-entropy terms, ensuring the action predictor selects the correct action, with a sparsity regularization term that keeps the attention map controllable.
@@
 After training, the CTR attention models can be used independently of the autoencoder and action predictor.
-It is noted that the CTR attention network can be used plug-and-play with the self-supervised attention network from [6], which has been shown to improve the rate of convergence and performance of deep RL methods in Atari games, while additionally offering task-relevant and contextualized feature selection.
+It is noted that the CTR attention network can be used plug-and-play with the self-supervised attention network from [wu2021self], which has been shown to improve the rate of convergence and performance of deep RL methods in Atari games, while additionally offering task-relevant and contextualized feature selection.
 This will therefore be investigated in Sec.
@@
 This approach would extract all features that are evidence of an action and could be more robust for applications in RL.
-Forward action prediction, equivalent to behavioral cloning, is used in this study as it leads the attention map to extract features that likely caused the player’s decision.
+However, we use forward action prediction as it directly extracts features that preceded and likely influenced the player’s decision, aligning with our goal to reveal decision-relevant attention patterns.
 2.3 Temporally-Integrated Overt Attention (TIOA) Network
 In order to analyze whether the CTR network can reveal human internal attention, a point of comparison is needed.
-Rather than using only the instantaneous gaze positions (eye-tracking data) from the Atari-HEAD dataset, an integration of recent gaze positions is preferable, as it more closely mimics the temporal integration of visual sensory data in humans [30].
+Rather than using only the instantaneous gaze positions (eye-tracking data) from the Atari-HEAD dataset, an integration of recent gaze positions is preferable, as it more closely mimics the temporal integration of visual sensory data in humans [fairhall2014temporal].
 This is relevant for comparison in accordance to our definition in Sec.
@@
 Figure 4: Action predictor accuracies over unseen attention-masked states for human (CTR and TIOA) and agent (only CTR) gameplay at different attention activation rates/ sparsities.
+Dashed lines indicate the minimum activation rate where 97% of the maximum action prediction accuracy is reached.
+Figure 5: Relative percentage decrease in action prediction accuracy from maximum comparing human CTR vs.
+agent CTR across different sparsity levels.
+Mean and 95% confidence interval across six games.
+Agent attention shows approximately twice the relative accuracy decrease compared to human attention, supporting that agents have more distributed attention.
 In this section, we evaluate the action prediction accuracy for both human and agent gameplay under full observation and under attention masking.
@@
 where $\tau_{k}$ is the threshold such that $r_{\Gamma}$ fraction of features are activated.
-We observe that accuracies for the human actions are higher than in the original Atari-HEAD study [17], likely as we are predicting based on frame-stacked states.
+We observe that accuracies for the human actions are higher than in the original Atari-HEAD study [zhang2020atari], likely as we are predicting based on frame-stacked states.
 In total, four main findings can be observed.
@@
 This is the case for all games except Freeway.
-Secondly, the CTR attention maps are more efficient than the gaze data-based TIOA maps, achieving significantly higher accuracies with fewer activated features.
+Secondly, the CTR attention maps are significantly more efficient than the gaze data-based TIOA maps, achieving higher action prediction accuracies with fewer activated features.
+This finding is supported by one-sided t-tests; for example, at a 16% activation rate (as used in the RL experiments in Sec.
+5), where each game’s mean accuracy is treated as an independent sample, we observe that action prediction accuracy under the human CTR masks is significantly higher than under the TIOA masks ($p=0.0093$).
 This may be expected, as the CTR networks are optimized for this task.
@@
 The dashed lines in Fig.
-4 indicate the minimum activation rate where 95% of the maximum action prediction accuracy is reached.
-Thirdly, human actions can be accurately predicted with less features than agent actions.
-This agrees with the finding of Guo et al.
-[23] which was based only on gaze data.
-Humans focus on less objects than agents trained for long-term future reward, even if we include potential objects of covert attention in the periphery beyond the gaze.
-Lastly, despite the expected greater gameplay variety in the human replay memory (due to the involvement of four different human subjects), the prediction accuracies are similar to those of a replay memory of a single trained agent.
-The findings suggest that the CTR network correctly identifies contextually task-relevant features.
+4 indicate the minimum activation rate where 97% of the maximum action prediction accuracy is reached.
+Thirdly, human actions can be accurately predicted with fewer features than agent actions, consistent with the conclusion of Guo et al.
+[guo2021machine] from an analysis of human gaze data.
+Our results extend this finding to decision-relevant attention patterns extracted via behavioral cloning, capturing internal attention beyond the gaze peak.
+Fig.
+5 quantifies this difference: agent attention shows approximately twice the relative accuracy decrease from maximum compared to human attention, though with large confidence intervals reflecting game-dependent variation.
+While this supports Guo et al.’s finding, direct quantitative comparison is limited as their work compared agent saliency maps with human gaze peaks (single concentrated regions), whereas our CTR approach captures broader internal attention including covert attention and peripheral awareness, suggesting the human-agent attention breadth difference may be less pronounced as it appears from gaze-focused analyses.
+Lastly, despite the expected greater gameplay variety in the human replay memory (due to the involvement of four different human subjects), the prediction accuracies vary by game but are generally similar to those of a replay memory of a single trained agent.
+The findings suggest that the CTR network correctly identifies contextually decision-relevant features.
 Whether these maps plausibly resemble human internal attention patterns will be further analyzed quantitatively and qualitatively in Sec.
 4.
-4 Evaluation II: Analysis Revealed Attention Maps
+4 Evaluation II: Analysis of Revealed Attention Maps
 4.1 Visual Analysis of Attention Maps
@@
 CTR human attention maps more closely resemble the TIOA than the agent attention.
-They likely represent human interal attention patterns as they include parts of the TIOA map, objects nearby, and highly task-relevant objects (see annotations).
+They likely represent human internal attention patterns as they include parts of the TIOA map, objects nearby, and highly decision-relevant objects (see annotations).
 Figure 7: Smooth attention map comparison: Example scenes (unseen states) with TIOA and Human CTR attention map that have been evaluated at multiple sparsity levels, smoothed and averaged.
@@
 In Fig.
-5, a scene is presented for each game with overlays of the TIOA and CTR maps.
+6, a scene is presented for each game with overlays of the TIOA and CTR maps.
 The selected scenes are intended to represent typical characteristics of the TIOA and CTR maps.
 Animations are provided in the supplemental materials.
-To adjust for differences between human and agent attention differences identified in Sec.
-3 we base the chosen activation rates $r_{\Psi}$ off the 95% of action prediction accuracy threshold from 4, with a lower limit of $0.04$ and an upper limit of $0.16$ resulting in the values shown in Tab.
+To adjust for differences between human and agent attention patterns identified in Sec.
+3 we base the chosen activation rates $r_{\Psi}$ off the 97% of action prediction accuracy threshold from 4, with a lower limit of $0.04$ and an upper limit of $0.16$ resulting in the values shown in Tab.
 1.
@@
    Game    Agent    Human
-   Enduro    0.16    0.097
-   Freeway    0.096    0.04
-   MsPacman    0.16    0.077
-   Seaquest    0.143    0.04
-   SpaceInvaders    0.16    0.04
-   Riverraid    0.16    0.041
+   Enduro    0.065    0.078
+   Freeway    0.04    0.04
+   MsPacman    0.112    0.066
+   Seaquest    0.054    0.04
+   SpaceInvaders    0.059    0.04
+   Riverraid    0.064    0.048
 Table 1: CTR attention map activation rates $r_{\psi}$ for each game used in the analysis in Sec. 4.
 For the visualization of the CTR attention maps in Fig.
-5, the maps are binarized at a threshold of $0.5$, and then upscaled.
+6, the maps are binarized at a threshold of $0.5$, and then upscaled.
 The TIOA map probabilities are binarized at a threshold in the range $[0.005,0.01]$ depending on the game [^3] TIOA binarization thresholds are: $0.005$ (Freeway), $0.0075$ (Seaquest), $0.01$ (Enduro, MsPacman, SpaceInvaders, Riverraid)..
@@
 1 over all unseen states (validation dataset).
-It will be qualitatively analyzed for each game whether the TIOA and attention maps differ, and if the CTR attention maps plausibly resemble internal attention.
-Again, the TIOA maps serve as a point of comparison as no true label of the internal attention is available.
+It will be qualitatively analyzed for each game whether the TIOA and attention maps differ, and if the CTR attention maps plausibly resemble internal attention patterns.
+Again, the TIOA maps serve as a point of comparison as no true label for the decision-relevant features exists.
 Generally, the TIOA maps are expected to overlap better with the human CTR attention maps but not fully.
@@
 In Freeway, the agent consistently attends to most cars and the player’s avatar (a chicken crossing the road).
-In contrast, human CTR attention alternates between two distinct modes: when walking, the focus is mostly on the avatar (seen in 6), and when walking, attention shifts toward the road ahead and cars (seen in 5).
+In contrast, human CTR attention alternates between two distinct modes: when walking, the focus is mostly on the avatar (seen in 7), and when walking, attention shifts toward the road ahead and cars (seen in 6).
 The 2 modes can also be seen in quantiative analysis in the following Sec.
@@
 For MsPacman and Seaquest, the TIOA predicts human gaze primarily around the avatar character and a few opponents or collectibles.
-The human CTR attention map likely reflects a pattern of interal attention, especially in MsPacman, where the similarities are strongest.
+The human CTR attention map likely reflects a pattern of internal attention, especially in MsPacman, where the similarities are strongest.
 In the shown scene, the CTR attention map focuses on a relevant nearby ghost which is also looked at by the human avatar, but rather at the edge of focus area.
@@
 We observe that both human and agent CTR attention maps outperform the baseline of $1.0$.
-For all games, the human CTR maps reach a higher alignment score averaging at $3.26$ compared to the agent CTR maps, which reach $2.49$.
+Across all games, the human CTR maps achieve a higher average alignment score of $3.26$, while the agent CTR maps reach an average of $2.49$.
+To assess the statistical significance of this difference, we performed a one-sided t-test using the average alignment score for each game as an independent sample.
+The result indicates a statistically significant difference in favor of the human CTR maps, with a $p$-value of $0.00031$.
 The alignment scores vary across games, with the highest human-agent gap in Freeway and MsPacman, indicating that human attention patterns in these games are more closely aligned with TIOA predictions.
@@
 Interestingly, the alignment score per bin also spikes here as well, underlying that this pattern exhibits consistency with the TIOA predictions.
-Overall, these results demonstrate that human CTR attention maps consistently exhibit a strong similarity to the TIOA maps across all games, indicating that the CTR network can capture human internal attention patterns.
-5 Evaluation III: Human Internal Attention-Guided Reinforcement Learning
-Figure 9: Comparison of learning curves of the Dueling DDQN agent with PER for the plain agent, the human internal attention-guided agent (CTR), and the inverted attention agent.
+Overall, these results demonstrate that human CTR attention maps consistently exhibit a strong similarity to the TIOA maps across all games, indicating that the CTR network can reveal human internal attention patterns.
+While the CTR network extracts generalized patterns rather than exact momentary attention states, the combination of high action prediction accuracy under masking and statistically significant gaze alignment provides empirical support that these patterns reflect the features causally relevant to human decision-making in these games.
+5 Evaluation III: Human Attention-Guided Reinforcement Learning
+Figure 10: Comparison of learning curves of the Dueling DDQN agent with PER for the plain agent, the human attention-guided agent (CTR), the TIOA-guided agent, and the inverted attention agent.
 Three runs are averaged per agent type, after performing a rolling average of 200 steps per run.
 Shaded area indicates standard deviation.
-Overall, CTR-masked agents slightly outperform plain ones, showing stable learning.
-Achieving human scores in Atari games is a challenging benchmark for RL agents, especially under limited training steps.
-Even algorithms such as Rainbow Deep Q-Learning (DQN) [32] that combine several improvements over normal DQN fail to reach mean human-normalized performance within 10 million environment steps, while human-outperforming methods like Agent57 [33] are highly tailored for Atari.
+Overall, CTR-masked agents slightly outperform plain ones, showing stable learning, while TIOA-masked agents perform poorly.
+Achieving human-level scores in Atari games is a challenging benchmark for RL agents, especially under limited training steps.
+Even algorithms such as Rainbow Deep Q-Learning (DQN) [hessel2018rainbow] that combine several improvements over normal DQN fail to reach mean human-normalized performance within 10 million environment steps, while human-outperforming methods like Agent57 [badia2020agent57] are highly tailored for Atari.
 Guo et al.
-[23] found that RL agents (here, using Proximal Policy Optimization, PPO), when they fail, are more likely to neglect regions attended to by humans.
-We investigate whether the human internal attention maps can improve learning performance, by making RL agents focus on objects that are task-relevant for humans.
-This represents an alternative approach to imitation learning, as it aims to transfer not just actions but the underlying attentional strategies derived from those actions.
-As a strong baseline, we use DQN with the modifications of double DQN (DDQN) [34], dueling DQN [35], and proportional prioritized experience replay (PER) [36].
+[guo2021machine] found that RL agents (here, using Proximal Policy Optimization, PPO), when they fail, are more likely to neglect regions attended to by humans, and that agent attention becomes more similar to human gaze when reducing the discount factor from 0.99 to 0.9, suggesting that human attention patterns can guide RL agents toward more human-like behavior.
+We investigate whether the human attention patterns can improve learning performance, by making RL agents focus on objects that are decision-relevant for humans.
+This represents an alternative approach to imitation learning, as it aims to transfer not only actions but the underlying attentional strategies derived from those actions.
+As a strong baseline, we use DQN with the modifications of double DQN (DDQN) [van2016deep], dueling DQN [pmlr-v48-wangf16], and proportional prioritized experience replay (PER) [schaul2015prioritized].
 Dueling DQN improves learning stability by separating value and advantage streams, double DQN mitigates overestimation bias by using a separate target network, and PER allows the agent to learn more effectively from rare but significant events.
-These enhancements are implemented using a custom-modified version of the stable baselines 3 (SB3) [37] DQN implementation.
+These enhancements are implemented using a custom-modified version of the stable baselines 3 (SB3) [JMLR:v22:20-1364] DQN implementation.
 The policy network receives the feature representation $\mathcal{F}_{\mathrm{p}}$ as input.
-We evaluate three agent variants: (1) a plain agent using features from a pre-trained, frozen autoencoder ($\mathcal{F}_{\mathrm{p}}=\mathcal{F}$), (2) a human internal attention-guided agent, where features are masked by the attention map $\bm{\Psi}$ to focus on task-relevant objects ($\mathcal{F}_{\mathrm{p}}=\bm{\Psi}\odot\mathcal{F}$), and (3) an inverted attention agent, where features are masked by the inverted CTR map ($\mathcal{F}_{\mathrm{p}}=(1-\bm{\Psi})\odot\mathcal{F}$).
-The comparison to the inverted map is done similarly to [38], providing a control baseline to evaluate the usefulness of the attention maps, especially since selective attention does not always guarantee performance gains [39].
-For each agent (1-3), we perform three runs of different seeds.
+We evaluate four agent variants: (1) a plain agent using features from a pre-trained, frozen autoencoder ($\mathcal{F}_{\mathrm{p}}=\mathcal{F}$), (2) a human attention-guided agent, where features are masked by the CTR attention map $\bm{\Psi}$ to focus on decision-relevant objects ($\mathcal{F}_{\mathrm{p}}=\bm{\Psi}\odot\mathcal{F}$), with attention values below 0.1 set to zero to ensure no information leakage at low attention values, (3) a TIOA-guided agent, where features are masked by binarized TIOA gaze prediction maps ($\mathcal{F}_{\mathrm{p}}=\hat{\bm{\Gamma}}_{\mathrm{bin}}\odot\mathcal{F}$), and (4) an inverted attention agent, where features are masked by the inverted CTR map ($\mathcal{F}_{\mathrm{p}}=(1-\bm{\Psi})\odot\mathcal{F}$).
+The comparison to the inverted map is done similarly to [itaya2021visual], providing a control baseline to evaluate the usefulness of the attention maps, especially since selective attention does not always guarantee performance gains [yuezhang2018initial].
+The TIOA-guided agent allows us to assess whether CTR attention provides additional value beyond overt attention alone for feature-level masking.
+For each agent (1-4), we perform three runs of different seeds.
 For all experiments, we use $16\%$ average activation rate in the CTR attention maps (meaning $84\%$ average activation for the inverted maps).
+For the TIOA-guided agents, the binarization threshold is set to achieve the same $16\%$ average activation rate, enabling a fair comparison of feature selection ability at matched sparsity levels.
+We apply hard masking for both CTR and TIOA to directly compare their effectiveness at selecting relevant features, though other approaches to incorporate gaze information (e.g., as auxiliary input channels or through soft masking) are possible and may perform differently.
 We set the Dueling DDQN and PER hyperparameters to standard values commonly used in the literature.
@@
 Fig.
-9 shows the learning curves for the three agent variants, denoting the reward obtained during the actual training episodes.
-The human internal attention-guided agent outperforms the plain agent for Enduro, Freeway and MsPacman, achieves similar performance for SpaceInvaders and slightly worse for Seaquest and Riverraid.
+10 shows the learning curves for the four agent variants, denoting the reward obtained during the actual training episodes.
+The human CTR-guided agent outperforms the plain agent for Enduro, Freeway and MsPacman, achieves similar performance for SpaceInvaders and slightly worse for Seaquest and Riverraid.
 The inverted attention agent performs worse in most games, especially in Seaquest where almost no learning progress is made.
@@
 The CTR-masked agents also appear to learn quite stably, especially in Enduro and MsPacman.
-Figure 10: Comparison of episodic scores of the Dueling DDQN agent with PER for the plain agent, the human internal attention-guided agent (CTR), and the inverted attention agent, including the random agent score.
+The TIOA-masked agents perform poorly across all games, worse than the inverted CTR agents except for Seaquest.
+Figure 11: Comparison of episodic scores of the Dueling DDQN agent with PER for the plain agent, the human CTR-masked agent, the TIOA-masked agent, and the inverted attention agent, including the random agent score.
 Three runs are averaged per agent type.
-Shaded area indicates standard deviation.
-Overall, CTR-masked agents consistently achieve higher episodic scores compared to plain agents.
+CTR-masked agents in average achieve higher episodic scores, while TIOA-masked agents perform poorly.
 Fig.
-10 shows the average episodic scores for the three agent variants after training.
-Here, the agents are evaluated under an exploration rate of $\epsilon=0.01$ for 150 episodes (50 per agent of different seeds), meaning they exploit their learned policies while still choosing a random action in $1\%$ of the cases.
+11 shows the average episodic scores for the four agent variants after training.
+Here, the agents are evaluated under an exploration rate of $\epsilon=0.01$ for 150 episodes (50 per seed) per agent type, meaning they exploit their learned policies while still choosing a random action in $1\%$ of the cases.
 An overall average relative score is determined by
 $\bar{R}^{(m)}=\frac{1}{n_{\mathrm{g}}}\sum_{g=1}^{n_{\mathrm{g}}}\frac{S_{g}^{(m)}-S_{g}^{(\mathrm{rand})}}{S_{g}^{(\mathrm{plain})}-S_{g}^{(\mathrm{rand})}}$ (17)
-is determined using random scores [^5] Random agent scores: Enduro: 0.0, Freeway: 0.0, MsPacman: 307.3, Seaquest: 68.4, SpaceInvaders: 148.0, Riverraid: 1338.5, taken from [33] and plain scores as the baselines.
+is determined using random scores [^5] Random agent scores: Enduro: 0.0, Freeway: 0.0, MsPacman: 307.3, Seaquest: 68.4, SpaceInvaders: 148.0, Riverraid: 1338.5, taken from [badia2020agent57] and plain scores as the baselines.
 The CTR-masked agent achieves highest episodic scores for MsPacman, Riverraid, and SpaceInvaders.
-It clearly outperforms the inverted attention agent in all games except Freeway, where all three agent types perform similarly.
-For the average relative score, the CTR-masked agents achieve $1.10$ whereas the inverted attention agents underperform with a score of $0.68$.
-The human internal attention-guided agents therefore reach $10\%$ higher scores than the baseline Dueling DDQN agent with PER.
-This shows that the derived attention maps can slightly improve RL agents performance, limiting the policy network’s available feature space to the $16\%$ predicted to be most relevant for humans.
+It clearly outperforms the inverted attention agent in all games except Freeway, where the plain, CTR-masked, and inverted attention agents perform similarly, while TIOA-masked agents perform worse.
+The TIOA-masked agents perform poorly across all games, even outperformed by inverterted CTR-masked agent except in Seaquest.
+For the average relative score, the CTR-masked agents achieve $1.10$, while the inverted attention agents underperform with a score of $0.68$, and the TIOA-masked agents achieve $0.52$.
+Thus, the human CTR-guided agents reach $10\%$ higher scores than the baseline Dueling DDQN agent with PER.
+To assess statistical significance, we conducted one-sided t-tests using each game’s average relative score as an independent sample per model.
+The CTR-masked agents significantly outperformed both the TIOA-masked agents ($p=0.0024$) and the inverted attention agents ($p=0.0162$).
+While the comparison with the plain, baseline agents did not reach statistical significance ($p=0.238$), the CTR-masked agents still achieved 10% higher scores on average across the tested games.
+Overall, although the performance improvement over baseline is not statistically significant across all games, these findings indicate a positive trend suggesting that CTR attention masks may improve RL agent performance.
+The lower performance of TIOA-masked agents suggests that binarizing gaze maps does not effectively select decision-relevant feature channels at the encoded level.
+In contrast, CTR attention learns to select feature channels that capture decision-relevant patterns, demonstrating its value beyond overt attention.
+We note that this comparison evaluates hard feature selection at matched sparsity; other approaches to incorporate gaze information may perform differently.
 6 Conclusions
-Our study presents a novel method to derive human internal attention patterns from gameplay data alone, leveraging offline attention methods from reinforcement learning (RL).
+Our study presents a novel method to derive human attention patterns from gameplay data alone, leveraging offline attention methods from reinforcement learning (RL).
 By introducing the contextualized, task-relevant (CTR) attention network, we generate attention maps from human and agent gameplay, compared against a temporally-integrated overt attention (TIOA) model based on eye-tracking data.
 After qualitative and quantitative evaluation, these maps are applied for attention-guided reinforcement learning.
-Evaluation across six Atari games demonstrates that human CTR attention maps exhibit higher alignment towrad the TIOA maps compared to those generated for agents.
-Overall, the human CTR maps capture key regions of the TIOA maps, as well as adjacent objects and highly task-relevant features, providing evidence that the CTR network reveals internal human attention patterns.
-Applied to RL, CTR-masked agents achieved 10% higher scores than baseline using only 16% of the feature space, demonstrating improved learning efficiency and performance, offering an alternative to imitation learning approaches.
+Evaluation across six Atari games demonstrates that human CTR attention maps identify features relevant for action prediction, and reach statistically significantly higher alignment toward the TIOA maps compared to those generated for agents.
+Overall, the human CTR maps capture key regions of the TIOA maps, as well as adjacent objects and highly decision-relevant features, providing evidence that the CTR network reveals human internal attention patterns.
+Applied to RL, CTR-masked agents achieved 10% higher scores than baseline over the tested games, using only 16% of the feature space.
+This hints at improved learning efficiency and performance and offers an alternative to imitation learning approaches.
+They further statistically significantly outperform TIOA-masked agents and inverted CTR-masked agents, reaching a $2.1$ and $1.6$ times higher episodic score respectively.
 Limitations include the need for larger human gameplay datasets with eye-tracking to improve accuracy and reduce noise in the CTR maps.
@@
 Methods could further be extended to datasets from autonomous driving or robotics.
-Our results represent a first step toward predicting human internal attention from interactions alone, without relying on additional data such as brain activity.
+Our results represent a first step toward predicting human internal attention patterns from interactions alone, without relying on additional data such as brain activity.
 This approach can deepen understanding of human-agent attention differences and contribute to human-inspired autonomous agents.
-Future work could explore more advanced strategies for human internal attention-guided learning, aiding the development of explainable, interpretable, and robust autonomous systems.
+Future work could explore more advanced strategies for human attention-guided learning, aiding the development of explainable, interpretable, and robust autonomous systems.
 […]
-Supplementary information
-We provide supplementary animations of the CTR and TIOA attention maps, namely the binarized maps shown in Fig.
-5, the smoothed maps shown in Fig.
-6, as well as the maps binarized at average activation of 16% as used in the RL experiments in Sec.
+Supplementary information We provide supplementary animations of the CTR and TIOA attention maps, namely the binarized maps shown in Fig.
+6, the smoothed maps shown in Fig.
+7, as well as the maps binarized at average activation of 16% as used in the RL experiments in Sec.
 5.
 Declarations
-The authors did not receive support from any organization for the submitted work.
+Funding: The authors did not receive support from any organization for the submitted work.
+Competing Interests: The authors declare no competing interests.